OpenStudy (anonymous):

A right triangle has a hypotenuse that measures 73 mm and one side that measure 48 mm. What is the length of the remaining side of the triangle?

OpenStudy (anonymous):

@iGreen A.55 mm B.60 mm C.65 mm D.75 mm

OpenStudy (kidrah69):

\[a^2+b^2=c^2\] plug the numbers

OpenStudy (kidrah69):

you know that the hypotenuse is c^2 so it will be 73^2

OpenStudy (kidrah69):

\[a^2+b^2=73^2\]

OpenStudy (kidrah69):

plug the other number in

OpenStudy (anonymous):

b=48^2?

OpenStudy (kidrah69):

looks right so you have \[a^2+48^2=73^2\] now solve this do you know how?

OpenStudy (nurali):

73^2 = a^2 + 48^2

OpenStudy (nurali):

a^2=73^2-48^2

OpenStudy (nurali):

a=sqrt(73^2-48^2)

OpenStudy (nurali):

a=?